OpenStudy (anonymous):

Identify the GCF of 12a4b3 + 8a3b2

OpenStudy (anonymous):

Greatest common factor is easiest to do when you simplify everything. \[12a^4b^3=2*2*3*a^4b^3\] \[8a^3b^2=2*2*2*a^3*b^2\] If you take the quotient of the two, you'll find all the ones that are able to cancel will make up your GCF \[\frac{2*2*3*a^4b^3}{2*2*2*a^3*b^2}\] Notice how the common terms that cancel are \[2*2*a^3*b^2\] If you write the factor out, you'll see it like this: \[2*2*a^3*b^2(3*ab+2)\]